OpenStudy (darkbluechocobo):

Find all the possible rational zeros of the following polynomial: f(x) = 3x^3 − 20x^2 + 33x − 9

OpenStudy (amistre64):

define descartes rule of sign for me

OpenStudy (darkbluechocobo):

@amistre64 basically it is the change from number to number if 5x -5x that is one change

OpenStudy (amistre64):

smartscores are out, looks lke another no notif day on openstudy ....

OpenStudy (amistre64):

but yes, we count the number of changes in 'sign'. postive roots are when we have a postivie x, negative with a negative x so i just use x=1 and x=-1 to represent them. notice that (-1)^odd = -1 so that effectively just reverses the sign on odd powers

OpenStudy (darkbluechocobo):

Yes but what changes when you want to find rational zeros?

OpenStudy (amistre64):

ah misread it ... rational zeros come from a pool of options that are constrcuted from the 'leading' coefficient and the constant tem

OpenStudy (amistre64):

if it has a rational zero, then it must be of the form:\[\pm\frac{first.coeff.factors}{last.term.factors}\]

OpenStudy (darkbluechocobo):

soh we have to find the factors for 3x^3 and 9?

OpenStudy (amistre64):

well, 3 and 9 yes

OpenStudy (darkbluechocobo):

well 3 and 1 for 3 and 3 and 3 for 9?

OpenStudy (amistre64):

\[\pm\frac{1,3}{1,3,9}\] what are all the possible combinations for this? this creates a pool of options for the only possible ratioanl roots that it can have

OpenStudy (darkbluechocobo):

well 1*9 and 3*3 is that what you mean?

OpenStudy (amistre64):

1/1, 1/3, 1/9 3/1, 3/3, 3/9 simplify and a set has no duplicate elements 1/1 = 3/3 so 1 is a solution

OpenStudy (amistre64):

possible solution

OpenStudy (amistre64):

1/3 = 3/9 so 1/3 is the possible solution

OpenStudy (darkbluechocobo):

3/1 which is 3 is 3 a possible solution?

OpenStudy (amistre64):

yes, it has no duplication so there is no other element to compare it to is all

OpenStudy (amistre64):

1/1, 1/3, 1/9 3/1, 3/3, 3/9 1, 1/3, 1/9 3, 1, 1/3 1, 1/3, 1/9, 3 +- of course

OpenStudy (darkbluechocobo):

Thank you for the helpa

OpenStudy (amistre64):

yep